Find the rms value, average value and form factor of resultant current in a circuit with simultaneous dc supply of 5A and ac supply with magnitude of 5A.
1
Expert's answer
2020-12-17T07:28:18-0500

1) "I_{rms}=\sqrt{\frac{5^2+(0.707\cdot5)^2}{2}}=4.3(A)"


2) "I_{avr}=\frac{5+0.637\cdot5}{2}=4.1(a)"


3) Form Factor"=\frac{I_{rms}}{I_{avr}}=\frac{4.3}{4.1}=1.05"